OpenStudy (anonymous):

Simplify the expression 4 over quantity of 3 x to the negative second power.

OpenStudy (anonymous):

\[\LARGE \frac{4}{3\cdot x^{-2}}\] are you asking this??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i dont understand it at all

OpenStudy (anonymous):

\[\LARGE \frac{4}{3}\cdot x^2\] just put x up and the power gets positive... I don't see nothing else to simplify !!.. O_O ... REMEMBER THIS : \[\LARGE \frac1x=x^{-1}\] and \[\LARGE x^{-1}=\frac1x\]